A sled is moving at a constant speed down a surface inclined at 45.0° with the horizontal and travels 30 meters in 4 seconds. Calculate the vertical velocity of the sled at the end of 4 seconds
1
Expert's answer
2013-01-18T09:02:30-0500
As speed is constant, V = d / t.
The vertical velocity Vx = V * cosA = d * cosA / t.
As d = 30 m and t = 4 s, V = 30 / 4 = 7.5 m/s and A = 45.0°,getting:
Vx = 30 * 1 / (4 * sqrt(2)) = 7.5 / sqrt(2) =approx. 5.3033 m/s.